OpenStudy (anonymous):

Find the difference. 2x / 4xy - 6 / 2x a. x - 3 / y b. x - 6y / 2xy c. x - 3 / 2xy - x d. -5 / 2 Help please.

OpenStudy (anonymous):

@elementwielder

OpenStudy (anonymous):

\[\frac{ 2x }{ 4xy }- \frac{ 6 }{ 2x }\]

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

simplify the individual terms first. 2x/(4xy) = 2/4 * x/x * 1/y = ? 6/(2x) = 6/2 * 1/x = ?

OpenStudy (anonymous):

It is B correct?

OpenStudy (anonymous):

simplify the equation so it is this\[\frac{ 1 }{ 2y }-\frac{ 3 }{ x }\]

OpenStudy (anonymous):

B is not correct because where would you have drawn the extra "y" from?

OpenStudy (anonymous):

so we automatically rule that out

OpenStudy (anonymous):

B is right.

OpenStudy (anonymous):

how so?

OpenStudy (anonymous):

(9/3x) - (4/2x^2) = (3/x) - (2/x^2) = 3x/x^2 - 2/x^2 = (3x - 2)/x^2

OpenStudy (anonymous):

wrong one xD

OpenStudy (anonymous):

what lol? which one is wrong?

OpenStudy (anonymous):

x/4xy - 6/2x = x/2xy - (6/2x)(y/y) = x/2xy - 6y/2xy = (x-6y)/2xy

OpenStudy (anonymous):

its B

OpenStudy (anonymous):

im wondering why in your second step you multiplied y/y to 6/2x

hartnn (hartnn):

to make the denominator common as xy and yes, its B

hartnn (hartnn):

as 2xy

OpenStudy (anonymous):

oh true... i was thinking they were equal not as on one side of the equation

OpenStudy (anonymous):

my bad